OpenStudy (anonymous):

help me out with Calculus! The angle between two differentiable curves at a point of intersection is the angle between the curves' tangent lines at this point. Find the angle between the curves in (a) and (b) at their points of intersection. (a) y= 1-x^2, y=x^2 (two points of intersection) (b) y=x^2, x=y^3 (two points of intersection)

OpenStudy (experimentx):

1) solve and find the points of intersection first. http://www.wolframalpha.com/input/?i=y%3D+1-x^2%2C+y%3Dx^2+ 2) find the tangents, (dy/dx) at that point gives the slope of tangent 3) find the angle ..

OpenStudy (anonymous):

can you please explain more?

OpenStudy (experimentx):

which step??

OpenStudy (anonymous):

2nd and 3rd

OpenStudy (experimentx):

find dy/dx of both, y= 1-x^2, y=x^2 dy/dx = -2x, dy/dx = 2x you know the values of x and y from step 1) right?? put those values, on above, you will have two slopes of tangents, m1 and m3 You know this formula right, \[ \tan(\theta) = \frac{m1 - m2}{1 + m1*m2}\] put those values of m1 and m2 and find the angle. And since it's difference, always take the positive value ... or negative won't affect it either.

OpenStudy (anonymous):

there are two points of intersection, so im not sure how i can only get two slopes of tangents

OpenStudy (anonymous):

can someone explain this

OpenStudy (experimentx):

well, find the slopes of tangent at one point of intersection... there you will have the angle between two curves. Then move to the other point of intersection, there you will have another angle of intersection. Good luck ...!!! it's going to lot of work!!!

OpenStudy (anonymous):

so im supposed to have 4 angles for part a?....

OpenStudy (anonymous):

can you teach me part a step by step and ill do part b.

OpenStudy (amistre64):

the points of intersection are when y = y

OpenStudy (amistre64):

y= 1-x^2, y=x^2 y = y 1-x^2 = x^2 and solve for x

OpenStudy (anonymous):

yeah i got the answer for that part, and then?:)

OpenStudy (amistre64):

what are the points of intersection? we will need them

OpenStudy (amistre64):

in calculus we learn that the derivative of an equation defines the slope of the tangent line at any given point

OpenStudy (anonymous):

\[(-1/\sqrt{2} , 1/2) , (1/\sqrt{2} ,1/2)\]

OpenStudy (amistre64):

goo, the x parts are all we will need inthis one since thats all we have to work with in the equations

OpenStudy (amistre64):

good, not goo lol

OpenStudy (anonymous):

haha

OpenStudy (amistre64):

what are our derivatives?

OpenStudy (anonymous):

dy/dx = -2x and dy/dx=2x

OpenStudy (amistre64):

good, plug in x = 1/sqrt(2) into them to find the 2 slopes we need to compare

OpenStudy (anonymous):

\[-\sqrt{2} and \sqrt{2}\]

OpenStudy (amistre64):

good. now all we need to do is determine the angle between them

OpenStudy (amistre64):